*assume after tax and DTS*

Retro Inc. has 1,100 bonds outstanding that are selling for $992 each. The bonds carry a 6.0 percent coupon, pay interest semi-annually, and mature in 7.5 years. The company also has 9,500 shares of 5% preferred stock at a market price of $40 per share. This month, the company paid an annual dividend in the amount of $1.20 per share. The dividend growth rate is 5.0 percent. The common stock is priced at $30 a share and there are 34,500 shares outstanding. The company is considering a project that is equally as risky as the overall company. This project has initial costs of $630,000 and operating cash flows of $80,000 a year for the next 10 years and salvage value of $20,000 at the end of 10 years. The net working capital (NWC) is expected to increase by $10,000 a year until the end of the project life. All the NWCs will be recovered when the project is completed. The project will be depreciated straight-line to zero over the project’s 10-year life. The tax rate is 21%.

  1. What is Retro’s weighted average cost of capital?

  1. What is the net present value (NPV) of this project? Should you accept the project? Explain why.
  1. What is the internal rate of return (IRR) of this project? Should you accept the project if you apply the IRR decision rule?
Answer #1

A. Yield of the bonds can be calculated in Excel as =RATE(15,30,-992,1000,0). This includes 15 payments of $30 each, with PV being 992 and future value being 1000. This is equal to 3.067% for 6 month period, or annual yield being 3.067*2 = 6.13%

Cost of debt = 6.13%*(1-tax rate) = 6.13%*(1-34%) = 4.05%

Cost of preferred equity = 5%

Price of stock = next year dividend / (cost of equity-growth rate)

So 30=1.2*(1+5%) / (cost of equity-5%), which means cost of equity = 9.2%

Weight of bonds in capital structure = (1100*992)/(1100*992+9500*40+30*34500) = 0.4354

Weight of preferred equity in capital structure = (9500*40)/(1100*992+9500*40+30*34500) = 0.1516

Weight of common equity in capital structure = (30*34500)/(1100*992+9500*40+30*34500) = 0.4130

WACC = 0.4354*4.05%+0.1516*5%+0.4130*9.2% = 6.32%

B. Depreciation = 630,000/10 = 63,000

Annual post-tax profit = (150,000-63,000)*(1-34%) = 57,420

Annual cashflow = 57,420+63,000=120,420

NPV = -630,000 + 120,420/(1+6.32%)^1 + 120,420/(1+6.32%)^2 + 120,420/(1+6.32%)^3 + ... + 120,420/(1+6.32%)^9 + 120,420/(1+6.32%)^10 + 10,000*(1-34%)/(1+6.32%)^10

= $246,630

As NPV is positive, the project should be accepted.